A balloon of mass is descending at a constant acceleration alpha .When a mass m is released from a balloon it starts rising with the same acceleration alpha.Assuming that it's value does not change,what is the value of m?

Grade:11

1 Answers

Arun
5 years ago
 

Let the force due to buoyancy be B.

In the first case the equation will be

Mg - B = Ms

In the second case the equation will be

B - (M-m)g = (M-m)s

Since the volume is unchanged, the force of Buoyancy remains unchanged as same amount of air is displaced even now and the new mass is (M-m)

Adding the two equations we get,

mg = (2M-m)s

or m(g+s)=2Ms

or m=2Ms/(g+s)